A girl has a collection of 425 coins consisting of pennies, nickels and dimes. she has 50 more nickels than pennies and 25 more dimes than nickels. how many coins of each kind does she have?
Answer by scott8148(6628)

p + n + d = 425

n = p + 50

d = n + 25 ___ d = (p + 50) + 25 ___ d = p + 75

p + (p + 50) + (p + 75) = 425

3p + 125 = 425

p = 100

n = 150

d = 175